OpenStudy (anonymous):

Giving medals!! What is the slope of the line with equation y = –x? A. 1 B. the slope is undefined C. -1 D. 0

OpenStudy (anonymous):

C

OpenStudy (anonymous):

You sure? Thanks!

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

One more question http://static.k12.com/calms_media/media/1299500_1300000/1299536/1/ec6e9876c2c49050325cab78f99a4e46e94b1d88/143685.jpg A.2x – y = 3 B.2x + y = –3 C.x – 2y = 6 D.x + 2y = –6